# Federated Learning with Dynamic Client Arrival and Departure: Convergence and Rapid Adaptation via Initial Model Construction

### v2

@@ Abstract
Most federated learning (FL) approaches assume a fixed client set. However, real-world scenarios often involve clients dynamically joining or leaving the system based on their needs or interest in specific tasks. This dynamic setting introduces unique challenges: (1) the optimization objective evolves with the active client set, unlike traditional FL with a static objective; and (2) the current global model may no longer serve as an effective initialization for subsequent rounds, potentially hindering adaptation. To address these challenges, we first provide a convergence analysis under a non-convex loss with a dynamic client set, accounting for factors such as gradient noise, local training iterations, and data heterogeneity. Building on this analysis, we propose a model initialization algorithm that enables rapid adaptation to new client sets whenever clients join or leave the system. Our key idea is to compute a weighted average of previous global models, guided by gradient similarity, to prioritize models trained on data distributions that closely align with the current client set, thereby accelerating recovery from distribution shifts. This plug-and-play algorithm is designed to integrate seamlessly with existing FL methods, offering broad applicability in practice. Experimental results on diverse datasets including both image and text domains, varied label distributions, and multiple FL algorithms demonstrate the effectiveness of the proposed approach across a range of scenarios.
1 Introduction
Federated learning (FL) is a decentralized machine learning paradigm that facilitates collaborative model training across multiple clients, such as smartphones and Internet of Things (IoT) clients, without exchanging individual data. Instead of transmitting raw data to the central server, each client performs local training using its proprietary data, sending only model updates to the server. These updates are then aggregated to refine the global model. In conventional FL frameworks, the cohort of clients engaged in training is typically static, implying the objective function is also fixed.
Motivation: However, in practical FL systems, the dynamic nature of client arrival and departure presents significant challenges to maintaining a robust and accurate model. For instance, clients may lose interest when their local data no longer aligns with the central task, such as when a user’s application shifts from text predictions to image recognition. On the other hand, clients may join when their current objectives align with those of other clients, such as when multiple users are working with similar data types or models addressing the same problem, like medical institutions collaborating on disease detection models. Additional factors, such as evolving privacy policies, or shifts in data-sharing preferences, further exacerbate these arrival and departure fluctuations, complicating the overall learning process.
Figure 1: Comparison between traditional FL with partial participation [1, 2, 3, 4, 5, 6] and our proposed dynamic FL framework. Traditional FL approaches assume a fixed client set and a static optimization goal, even when some clients are intermittently available. In contrast, our framework handles dynamic client arrivals and departures, where the optimization goal evolves to minimize the loss for the current client set in each session. A key innovation of our approach is the construction of an adaptive initial model for each session, leveraging a weighted averaging strategy guided by gradient similarity. This initial model prioritizes contributions from clients whose data characteristics align closely with the current set, facilitating efficient adaptation and mitigating the impact of client dynamics. Traditional FL with full client availability is treated as a special case of the static, partial-participation setting in this comparison.
Key Challenges: This dynamic FL setting introduces new challenges that are not present in traditional static FL scenarios with a fixed objective function. Specifically, the arrival and departure of clients dynamically alter the FL system’s objective, as the model must adapt to the current client set and their associated tasks. For instance, if a client contributing unique data classes withdraws, the model no longer needs to classify those classes, fundamentally altering the training objective. Conversely, the addition of clients with previously unrepresented classes necessitates model adaptation to incorporate the updated classification task. Thus, the core challenge is not merely preserving the diversity of the training set but dynamically adjusting the model to align with the evolving tasks defined by the current clients in the system. To address these shifting objectives, FL systems must be designed with the ability to rapidly adapt to the current client set, ensuring the model remains relevant, stable, and effective as the landscape of participating clients evolves.
Illustrative Examples: Figure 1 illustrates our system model, emphasizing its departure from traditional FL with a fixed client set. In conventional FL, although some clients may temporarily become inactive, the overall optimization goal remains static, aiming to minimize the aggregated loss across all clients and converge toward a single global model $\bm{w}^{*}$. Prior works typically focus on designing client sampling strategies (e.g., [6]) or modifying aggregation weights (e.g., [7]) to minimize this static global loss, without accounting for evolving client sets and objectives. In contrast, our approach introduces a session-based framework that inherently supports dynamic client patterns, where the set of active clients evolves across sessions. Instead of assuming a fixed client set throughout training, our system explicitly models client arrivals and departures at the session level. Each session optimizes a session-specific model based on the current client set, ensuring adaptability to shifting data distributions. To enhance efficiency, an initial model is constructed at the start of each session using the last global model in each previous sessions, providing a better starting point than simply resuming from the last model. Within each session, our scheme also supports partial client involvement, allowing training to proceed even when only a subset of clients is active in each round. By jointly addressing dynamic client arrivals across sessions and partial client involvement within sessions, our framework offers a more flexible and realistic approach.
@@ 1 Introduction
Main Difference from Prior Work: Unlike existing continual learning paradigms that aim to prevent catastrophic forgetting [8, 9, 10], our work focuses on enabling fast adaptation to the current data distribution. This resembles a service provider aiming to build a global model that effectively serves the clients who are currently active, motivated by scenarios (e.g., mobile/edge environments) where clients frequently join and leave. From this perspective, once a client leaves, the priority shifts to optimizing performance for the remaining users rather than preserving accuracy for those who have departed. We thus adopt a dynamic, session-centric view, emphasizing responsiveness to the active client population. In Section 6, we include class incremental learning as one of our baselines and show that its performance is much worse under our objective.
@@
Contributions: Given the challenges introduced by the dynamic nature of client arrival and departure, this work makes the following key contributions to advance the field:
@@
  • We provide a novel theoretical analysis with a non-convex loss function, deriving an upper bound on the gradient norm to quantify performance under any pattern of client arrival and departure in dynamic FL. Our analysis considers a dynamic optimization objective, where each session aims to identify an optimized model that minimizes the loss function of the current clients, contrasting with existing literature that primarily addresses a static goal of minimizing the loss across all clients that have ever participated. The derived bound incorporates several critical factors influencing performance, including: (1) stochastic gradient noise arising from inherent randomness in local updates, (2) the number of local training iterations, which affects convergence behavior, and (3) the non-IIDness of data distribution. This comprehensive analysis offers deeper insights into the trade-offs involved in model training under dynamic client arrivals and departures.
@@
  • Building on our analysis, we propose a plug-and-play algorithm designed to address dynamic optimization goals by constructing an effective initial model for each training session. The algorithm seamlessly integrates with existing FL methods (e.g., FedAvg [11], FedProx [12], SCAFFOLD [13], and FedACG [14] to facilitate rapid adaptation to a new client set in dynamic environments. It initializes the model as a weighted average of previous models, with weights determined by gradient similarity. By leveraging these similarities, the algorithm prioritizes models trained on clients whose data characteristics closely align with those of the current ones, thereby improving adaptation and mitigating performance drop caused by dynamic client activities.
  • Experimental evaluations on diverse datasets spanning both image and text domains demonstrate significant performance gains from our algorithm, particularly in scenarios with sporadic or moderately varying client patterns, highlighting the method’s effectiveness in real-world settings.
2 Related Works
@@
Federated learning (FL) has emerged as a prominent paradigm for distributed machine learning, enabling the collaborative training of models across decentralized data sources while preserving data privacy [11, 15, 12]. One of the foundational works introduced the Federated Averaging (FedAvg) algorithm, which remains a cornerstone of many FL systems [11]. Subsequent studies have explored various aspects of federated learning, including communication efficiency [16], robustness to adversarial attacks [17], and personalization strategies [18]. However, these approaches typically assume a fixed set of participating clients throughout the training process. Reviews of FL techniques often consider static client participation, where all clients are expected to remain available for the entire training duration [15, 12]. This assumption simplifies the modeling of convergence and performance but does not adequately capture real-world scenarios characterized by dynamic client pattern. Addressing the dynamic nature of client arrival and departure is a critical gap in the current literature, motivating the need for adaptive methods that can effectively handle the entry and exit of clients during the training process.
@@
To address the limitations of static client sets in federated learning, research on dynamic client selection and flexible participation has gained momentum, particularly in response to the challenges posed by varying client availability [19, 20, 21, 22, 23, 24, 25, 26, 27, 3]. These studies explore strategies such as optimizing client selection based on resource constraints, modeling participation patterns probabilistically, and employing adaptive algorithms to address the effects of non-IID data. The goal is to enhance overall model performance by strategically managing client participation during training, balancing computational efficiency, communication costs, and data representativeness. However, these works often assume a fixed client set, neglecting the dynamic nature of client arrival and departure. While [28] propose a flexible federated learning framework that allows for inactive clients, incomplete updates, or dynamic client participation, their analysis is limited to scenarios where the optimization goal changes only once, specifically when a single client joins during training. This restricts the applicability of their approach to more complex and realistic patterns of client pattern. Also, their analysis relied on more restrictive assumptions such as strong convexity and universally bounded gradient. There is a line of work in continual FL [29, 30] , where the goal is to build a model that preserves all previously learned knowledge by preventing catastrophic forgetting. In contrast, as previously mentioned, our work focuses on rapidly adapting to the current data distribution to maximize the satisfaction of clients currently active in the system. Given this difference in objectives, our approach significantly outperforms continual FL in scenarios involving dynamic client arrivals and departures, as we will see in Section 6.
3 Dynamic Federated Learning with Client Arrival and Departure
We consider a dynamic FL system where clients may join or leave the system based on their interest or task needs. Clients may join when the model aligns with their objectives, when they have sufficient new data to contribute, or when the system offers financial or computational incentives. Conversely, they may leave if the global model drifts from their needs, if they lack sufficient data, or if the model’s performance is not beneficial. Privacy or security concerns, such as adversarial threats or inadequate privacy guarantees, may also lead clients to leave. Additionally, clients may leave to allocate computational resources to other tasks or due to poor local model validation, while others may rejoin when they see improvements in these factors.
To mathematically characterize these scenarios, each round of FL training, denoted by $g\in\mathcal{G}=\{1,\dots,G\}$, involves a client set collected in $\mathcal{K}^{(g)}=\{1,\dots,K^{(g)}\}$. All clients are connected to a central server, with each client maintaining its own ML models tailored to specific tasks such as pattern recognition and natural language processing. These models are trained locally on client-specific data, allowing the system to leverage diverse data sources while preserving user privacy. At global round $g$, every client $k\in\mathcal{K}^{(g)}$ is linked with a dataset $\mathcal{D}_{k}^{(g)}$, which comprises $D_{k}^{(g)}\triangleq|\mathcal{D}_{k}^{(g)}|$ data samples. Each data sample $d\in\mathcal{D}_{k}^{(g)}$ includes a feature vector $d$ and its corresponding label. A loss function $f(\bm{w},d)$ evaluates the performance of parameter $\bm{w}\in\mathbb{R}^{M}$ where $M$ represents the model’s dimensionality for each data sample $d\in\mathcal{D}_{k}^{(g)}$. For a specific $\bm{w}$, during the $g$-th global round, the local loss at client $k$ is defined as $F_{k}^{(g)}(\bm{w})\triangleq\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{f(\bm{w},d)}{D_{k}^{(g)}}.$ Then, the global loss of the ML model is given by $F^{(k)}(\bm{w})\triangleq\frac{1}{D^{(g)}}\sum_{k\in\mathcal{K}^{(g)}}D_{k}^{(g)}F_{k}^{(g)}(\bm{w})$ where $D^{(g)}\triangleq|\mathcal{D}^{(g)}|$ is the cardinality of $\mathcal{D}^{(g)}$, the collection of data samples across clients. The main objective of training is to minimize the global loss function, which determines the model’s effectiveness in real-world downstream tasks on clients. However, dynamic client pattern, where clients may join or leave the system at any time, causes the global loss function to vary across rounds. Thus, the optimal global model parameters form a sequence $\{\bm{w}^{(g)\ast}\}_{g=1}^{G}$, where
$\bm{w}^{(g)\ast}=\arg\min_{\bm{w}\in\mathbb{R}^{M}}F^{(g)}(\bm{w},\mathcal{D}^{(g)}),\quad g\in\mathcal{G}.$ (1)
Given this dynamic objective, the training process in round $g$ is specifically designed to optimize the model for the data currently held by the participating clients, aiming to minimize the loss function $F^{(g)}(\bm{w},\mathcal{D}^{(g)})$. Unlike continual learning, which typically aims to retain performance on previously encountered data distributions, this training strategy focuses on optimizing for the data distribution specific to the current round $g$. In each training round, the server broadcasts the current global model $\bm{w}^{(g)}\in\mathbb{R}^{M}$ to all clients $\mathcal{K}^{(g)}$. Upon receiving the global model, each client $k\in\mathcal{K}^{(g)}$ updates the model using its local dataset $\mathcal{D}_{k}^{(g)}$ over $e_{k}^{(g)}$ steps of stochastic gradient descent (SGD). During the $h$-th SGD iteration, where $h\in\{0,\dots,e_{k}^{(g)}-1\}$, the update rule is given by $\bm{w}_{k}^{(g),h+1}=\bm{w}_{k}^{(g),h}-\eta^{(g)}\nabla F_{k}^{(g)}(\bm{w}_{k}^{(g),h},\mathcal{B}_{k}^{(g)}),$ where $\eta^{(g)}$ denotes the learning rate in the $g$-th round, and $\nabla F_{k}^{(g)}(\bm{w}_{k}^{(g),h},\mathcal{B}_{k}^{(g)})$ is the gradient of client $k$’s local loss function evaluated on the mini-batch dataset $\mathcal{B}_{k}^{(g)}\subset\mathcal{D}_{k}^{(g)}$. The mini-batch dataset is sampled from the local dataset $\mathcal{D}_{k}^{(g)}$ to compute the stochastic gradient. The initial model for local training is the current global model, i.e., $\bm{w}_{k}^{(g),0}=\bm{w}^{(g)}$, and the final local model after training is denoted as $\bm{w}_{k}^{(g),\mathsf{F}}$. After completing local training, clients send their final models $\bm{w}_{k}^{(g),\mathsf{F}}$ to the server. The server aggregates these models to update the global model as $\bm{w}^{(g+1)}=\sum_{k\in\mathcal{K}^{(g)}}(D_{k}^{(g)}\bm{w}_{k}^{(g),\mathsf{F}}/D^{(g)}).$ The server initiates the next training round by transmitting the updated global model $\bm{w}^{(g+1)}$ to the new client set $\mathcal{K}^{(g+1)}$. This set includes clients that joined before model aggregation and excludes those that left during the previous round. Unlike prior studies, which optimize for all potential clients, this work focuses on optimizing the model for the current set of participating clients in each round, effectively targeting $\bm{w}^{(g)\ast}$. This strategy aligns with the dynamic nature of client patterns, avoiding optimization for clients that may not rejoin. In this framework, existing clients must complete local training before leaving, while new clients can join anytime and participate in the next global round if they join after the model broadcast.
4 Theoretical Analysis
[…]

### v3

# Communication-Efficient Federated Learning under Dynamic Device Arrival and Departure: Convergence Analysis and Algorithm Design

@@ Abstract
Most federated learning (FL) approaches assume a fixed device set. However, real-world scenarios often involve devices dynamically joining or leaving the system, driven by, e.g., user mobility patterns or handovers across cell boundaries. This dynamic setting introduces unique challenges: (1) the optimization objective evolves with the active device set, unlike traditional FL’s static objective; and (2) the current global model may no longer serve as an effective initialization for subsequent rounds, potentially hindering adaptation, delaying convergence, and reducing resource efficiency. To address these challenges, we first provide a convergence analysis for FL under a dynamic device set, accounting for factors such as gradient noise, local training iterations, and data heterogeneity. Building on this analysis, we propose a model initialization algorithm that enables rapid adaptation whenever devices join or leave the network. Our key idea is to compute a weighted average of previous global models, guided by gradient similarity, to prioritize models trained on data distributions that closely align with the current device set, thereby accelerating recovery from distribution shifts in fewer training rounds. This plug-and-play algorithm is designed to integrate seamlessly with existing FL methods, offering broad applicability. Experiments demonstrate that our approach achieves convergence speedups typically an order of magnitude or more compared to baselines, which we show drastically reduces energy consumption to reach a target accuracy.
I Introduction
Federated learning (FL) facilitates collaborative model training across a set of devices connected to a central server. In FL, each device performs local training using its local data, and uploads model updates to the server periodically for aggregation. The aggregated models are then broadcasted across the devices to initiate the next round. FL has found many applications in resource-constrained environments, including the Internet of Things (IoT) [wang2019adaptive, abdulrahman2023management], autonomous driving [samarakoon2019distributed], and smart cities [xing2021federated].
In conventional FL frameworks [dinh2020federated, xiang2023federated, jhunjhunwala2022fedvarp, xu2020client], the cohort of edge devices engaged in training is typically assumed to be static, implying the objective function is fixed. Nevertheless, in practice, the assumption of a static device set rarely holds in wireless edge networks, where user mobility and fluctuating coverage patterns drive continuous system dynamics. Devices frequently traverse cell boundaries, triggering handovers that result in their departure from one federated system and potential entry into another. Similarly, time-varying channel conditions and finite battery capacities can force devices to drop out of the system entirely. Unlike static scenarios where a fixed device set simply “duty-cycles” due to intermittent connectivity, these mobility-driven events fundamentally alter the active topology of the network: as devices physically enter or exit the server’s coverage area, the system faces a structural evolution of the device set, necessitating a learning framework that can handle/track the available population in the network.
Figure 1: Comparison illustrating our proposed dynamic FL framework as a generalization of traditional FL. Traditional FL represents the single-session special case, limited to a fixed device set and a static optimization goal. Our framework generalizes this structure to accommodate dynamic device arrivals and departures across multiple sessions, allowing the optimization goal to evolve and minimize the loss for the currently active set of users.
Illustrative Example: Figure 1 illustrates our problem setting as a generalization of traditional FL to support dynamic device populations. Standard FL approaches typically handle temporary device unavailability by adjusting for partial participation, yet they assume the underlying set of users and the resulting optimization goal $\bm{w}^{*}$ remain fixed indefinitely. We generalize this by structuring model training into distinct sessions to accommodate population changes in the system (traditional FL corresponds to one single session). As shown in the bottom panel of Figure 1, we aim to retain the ability to handle partial device participation within a session while simultaneously adapting to device turnover between sessions. In doing so, the optimization goal will evolve from a single static target to a sequence of objectives specific to each session. By carefully accounting for device arrivals and departures, our framework establishes a sequence of session-specific optimization targets; however, efficiently tracking these shifting goals is crucial to preventing model stagnation as well as minimizing time to (re-)convergence and the associated communication overhead. This leads to our key research question:
Question: How can we design a plug-and-play method that (i) integrates with existing FL algorithms (e.g., FedProx [li2020federated], SCAFFOLD [karimireddy2020scaffold]) originally developed for static device sets within a single session and (ii) provides fast adaptation to the changing optimization goal $\bm{w}^{(s)\ast}$ as the system transitions between sessions $s$?
Realizing this adaptation is non-trivial as the system operates blindly with respect to future device dynamics: since the server has no prior knowledge of which devices will join or leave the system, it cannot pre-calculate the optimal starting point for the next session. Furthermore, simply carrying over the global model from the previous session is often suboptimal as significant shifts in the underlying data distribution can render the old model stale, forcing the algorithm to waste communication rounds merely “removing” outdated information. In resource-constrained wireless networks, this slow re-convergence is unacceptable, as it directly translates to excessive latency and higher energy consumption from executing more training rounds. The goal, therefore, is enabling fast global model re-convergence to reduce the resource overhead of each session.
@@ I Introduction
Difference from Continual Learning: Unlike existing continual learning paradigms that aim to prevent catastrophic forgetting [yoon2021federated], which we consider in Section VI as one of our baselines, our work focuses on enabling fast adaptation to the current data distribution. This resembles a service provider aiming to build a global model that effectively serves the users who are currently active, motivated by wireless edge scenarios where devices frequently join and leave a local network. From this perspective, once a device leaves, the priority shifts to optimizing performance for the remaining users rather than preserving accuracy for those who have departed. We thus adopt a dynamic, session-centric view, emphasizing responsiveness to the active device population.
Summary of Contributions
@@
To address the challenges introduced by dynamic device arrival and departure, this work makes the following key contributions:
@@
  • We conduct a novel theoretical analysis that analyzes the session-specific gradient norm under any pattern of device arrival and departure in dynamic FL. Our analysis considers a dynamic optimization objective, where each session aims to learn an optimized model that minimizes the loss function for the current devices. The bound incorporates critical factors influencing performance, including: (i) stochastic gradient noise arising from inherent randomness in local updates, (ii) the number of local training iterations, and (iii) the non-IID-ness of local data distributions.
@@
  • Building on our analysis, we propose a plug-and-play algorithm for dynamic optimization in FL, which constructs an effective initial model for each session. Aiming for low-complexity integration with existing FL methods (e.g., MOON [li2021modelmoon], FedProx [li2020federated], SCAFFOLD [karimireddy2020scaffold], FedACG [kim2024communication]), our algorithm initializes the model as a gradient similarity-weighted average of previous models, prioritizing models trained on devices whose data characteristics closely align with those of the current session. This approach accelerates model adaptation and mitigates costly performance drops caused by dynamic device participation.
  • We evaluate system performance using two distinct metrics: Time-to-Accuracy and Accumulated Accuracy Gain. Time-to-Accuracy measures the number of global rounds required to recover a specific target accuracy, which translates to adaptation latency and energy consumption as well. Our results show that our strategy reduces the rounds needed to reach a target accuracy typically by an order of magnitude or more compared to continual learning baselines. Through communication and computation modeling, we quantify the corresponding improvements in energy consumption as well. Complementing this, the Accumulated Accuracy Gain metric confirms superior sustained performance throughout the session, yielding total accumulated accuracy gains of up to 8x in challenging zero-overlap session scenarios.
II Related Works
@@
FL has emerged as a prominent paradigm for distributed learning, enabling the collaborative training of models across devices while preserving data privacy [mcmahan2017communication, li2020federated, nguyen2020fast]. One of the foundational works introduced the Federated Averaging (FedAvg) algorithm, which remains a cornerstone of many FL systems [mcmahan2017communication]. Subsequent studies have explored various aspects of FL, including enhancing resource efficiency [wang2021quantized, xing2021federated, luo2021cost, xu2023federated, chen2020joint], building robustness to adversarial attacks [wei2021low, yang2025communication], modeling energy consumption [yang2020energy], and more distributed (non-star) topologies [lin2021semi, xing2021federated, lim2021dynamic, luo2020hfel, yang2023decentralized, ye2022decentralized, chen2022decentralized, zhou2023decentralized]. However, these approaches typically assume a fixed set of participating devices throughout the training process, where devices are expected to remain available for the entire training duration [li2020federated]. This assumption simplifies the modeling of convergence and performance but does not adequately capture real-world scenarios characterized by dynamic device participation patterns.
@@
To address the limitations of static device sets in FL, research on dynamic device selection and flexible participation has gained momentum, particularly in response to the challenges posed by varying device availability [xu2020client, jhunjhunwala2022fedvarp, sun2021dynamic, ribero2022federated]. These studies explore strategies such as optimizing device selection based on resource constraints, modeling participation patterns probabilistically, and employing adaptive algorithms to address the effects of non-IID (independent and identically distributed) data. In essence, the goal of these studies is to enhance overall model performance by strategically managing device participation during training, balancing computational efficiency, communication costs, and data representativeness. However, these works often assume a fixed device set, neglecting the dynamic nature of device arrival and departure. While [ruan2021towards] propose a flexible FL framework that allows for inactive devices, incomplete model updates, or dynamic device participation, their analysis is limited to scenarios where the optimization goal changes only once, specifically when a single device joins during training. This restricts the applicability of their approach to more complex and realistic patterns of device availability. Also, their analysis relies on more restrictive assumptions such as strong convexity and universally bounded gradient. There is also a line of work in continual FL [dong2022federated], where the goal is to build a model that preserves all previously learned knowledge by preventing catastrophic forgetting. In contrast, as previously mentioned, our work focuses on rapidly adapting the global model to the current data distribution of devices. Given this difference in objectives, our approach significantly outperforms continual FL in scenarios involving dynamic device arrivals and departures, as we will show in Section VI.
III System Model and Theoretical Analysis
III-A System Model
We discretize the FL process into discrete sessions, indexed by $s\in\{1,\dots,S\}$. Each session consists of $T$ global training rounds, indexed by $t\in\{1,\dots,T\}$. To model device dynamics, $\mathcal{K}^{(s)}=\{1,\dots,K^{(s)}\}$ denotes the population of devices available in session $s$. Each device is connected to a central server and maintains its own ML model. Specifically, at session $s$, each device $k\in\mathcal{K}^{(s)}$ holds a dataset $\mathcal{D}_{k}^{(s)}$ of size $D_{k}^{(s)}\triangleq|\mathcal{D}_{k}^{(s)}|$, where each sample $d\in\mathcal{D}_{k}^{(s)}$ consists of a feature-label pair. A loss function $f(\bm{w},d)$ evaluates model parameters $\bm{w}\in\mathbb{R}^{M}$ over individual samples. The local loss at device $k$ during session $s$ is defined as follows:
$F_{k}^{(s)}(\bm{w})\triangleq\sum_{d\in\mathcal{D}_{k}^{(s)}}f(\bm{w},d)/D_{k}^{(s)},$ (1)
and the global loss function aggregates local losses over the current population:
$F^{(s)}(\bm{w})\triangleq\frac{1}{D^{(s)}}\sum_{k\in\mathcal{K}^{(s)}}D_{k}^{(s)}F_{k}^{(s)}(\bm{w}),$ (2)
where $D^{(s)}\triangleq|\mathcal{D}^{(s)}|$ denotes the total number of data samples across all available devices in session $s$.
In our scenario of interest, where devices may join or leave at any time, the FL training objective evolves per session. As a result, the optimal global models form a sequence $\{\bm{w}^{(s)\ast}\}_{s=1}^{S}$, with each $\bm{w}^{(s)\ast}$ corresponding to the current session’s data:
$\bm{w}^{(s)\ast}$ $=\arg\min_{\bm{w}\in\mathbb{R}^{M}}F^{(s)}(\bm{w},\mathcal{D}^{(s)}).$ (3)
This formulation is different from the traditional FL objectives (i.e., $\bm{w}^{\ast}=\arg\min_{\bm{w}\in\mathbb{R}^{M}}F(\bm{w},\mathcal{D})$), which seeks a single model $\bm{w}^{*}$ optimized over a fixed dataset $\mathcal{D}$. In contrast, our formulation targets a time-varying objective, optimizing $\bm{w}^{(s)\ast}$ based on the current/evolving data $\mathcal{D}^{(s)}$.
In our setting, each training round $t$ in session $s$ proceeds as follows. The server selects a set of participating devices $\mathcal{A}^{(s,t)}\subseteq\mathcal{K}^{(s)}$ and broadcasts the global model $\bm{w}^{(s,t)}\in\mathbb{R}^{M}$ to them. Each participating device $k$ then updates the model using its data $\mathcal{D}_{k}^{(s)}$ over $e_{k}^{(s,t)}$ steps of stochastic gradient descent (SGD). Specifically, across SGD iterations $h\in\{0,\dots,e_{k}^{(s,t)}-1\}$, the model update rule is as follows:
$\bm{w}_{k}^{(s,t),h+1}=\bm{w}_{k}^{(s,t),h}-\eta^{(s,t)}\nabla F_{k}^{(s)}(\bm{w}_{k}^{(s,t),h},\mathcal{B}_{k}^{(s,t)}),$ (4)
where $\eta^{(s,t)}$ is the learning rate and $\mathcal{B}_{k}^{(s,t)}\subset\mathcal{D}_{k}^{(s)}$ is a mini-batch. The local training starts from the global model $\bm{w}_{k}^{(s,t),0}=\bm{w}^{(s,t)}$ and results in a final local model $\bm{w}_{k}^{(s,t),\mathsf{F}}$. Devices send their final models to the server, which aggregates them as follows:
$\bm{w}^{(s,t+1)}=\sum_{k\in\mathcal{A}^{(s,t)}}\frac{D_{k}^{(s)}}{\sum_{j\in\mathcal{A}^{(s,t)}}D_{j}^{(s)}}\bm{w}_{k}^{(s,t),\mathsf{F}}.$ (5)
The updated global model $\bm{w}^{(s,t+1)}$ is then broadcast to devices for the next training round.
III-B Theoretical Analysis
To gain insights into how the FL algorithm performs in the complex system we consider, we first analyze its performance, which inspires the design of our proposed algorithm. We leverage the following two standard assumptions [wang2020tackling] to derive the analysis (henceforth, $\|\cdot\|$ denotes the two-norm).